OpenStudy (anonymous):

4. A shipping company is considering using this box to ship products.

OpenStudy (anonymous):

i guess surface area is SA=2*(lb+bh+lh) = 2*(12+6+8)=52 volume is vol = 4*3*2=24; clearly ratio is more than 2...